OpenStudy (anonymous):

Using implicit differentiation, find dy/dx. (e^y)cosx=8+sin(xy). I got [(e^y)sinx]/[(e^y)cosx-8cos(xy)y] but it was wrong.

OpenStudy (anonymous):

If I'm not mistaken, you should get:\[-e^y \sin x+y'e^y \cos x=0+ycos (xy)+y'x \cos (xy)\]solve for y'.

OpenStudy (anonymous):

did you use the chain rule and product rule?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

ok thank you